Problem 7-11 (Algorithmic)

Hart Manufacturing makes three products. Each product requires manufacturing operations in three departments: A, B, and C. The labor-hour requirements, by department, are as follows:

Department Product 1 Product 2 Product 3
A 1.60 3.10 2.10
B 2.20 1.20 2.70
C 0.35 0.35 0.35

During the next production period, the labor-hours available are 470 in department A, 370 in department B, and 70 in department C. The profit contributions per unit are $27 for product 1, $30 for product 2, and $32 for product 3.

  1. Formulate a linear programming model for maximizing total profit contribution. For those boxes in which you must enter subtractive or negative numbers use a minus sign. (Example: -300)

    Let Pi = units of product i produced
    Max P1 + P2 + P3
    s.t.
    P1 + P2 + 2.1P3
    2.2P1 + P2 + P3
    P1 + 0.35P2 + P3
    P1, P2, P3 ≥ 0
  2. Solve the linear program formulated in part (a). How much of each product should be produced, and what is the projected total profit contribution?

    P1 =  
    P2 =  
    P3 =  

    Profit = $ ______
  3. After evaluating the solution obtained in part (b), one of the production supervisors noted that production setup costs had not been taken into account. She noted that setup costs are $420 for product 1, $570 for product 2, and $620 for product 3. If the solution developed in part (b) is to be used, what is the total profit contribution after taking into account the setup costs?

    Profit = $ ______
  4. Management realized that the optimal product mix, taking setup costs into account, might be different from the one recommended in part (b). Formulate a mixed-integer linear program that takes setup costs into account. Management also stated that we should not consider making more than 185 units of product 1, 160 units of product 2, or 150 units of product 3. For those boxes in which you must enter subtractive or negative numbers use a minus sign. (Example: -300)

    Here introduce a 0-1 variable yi that is one if any quantity of product i is produced and zero otherwise.
    Max P1 + P2 + P3 + y1 + y2 + y3
    s.t.
    P1 + P2 + 2.1P3
    2.2P1 + P2 + P3
    P1 + 0.35P2 + P3
    P1 + y1
    P2 + y2
    P3 + y3
    P1, P2, P3 ≥ 0; y1, y2, y3 = 0, 1
  5. Solve the mixed-integer linear program formulated in part (d). How much of each product should be produced, and what is the projected total profit contribution? Compare this profit contribution to that obtained in part (c).

    P1 =  
    P2 =  
    P3 =  

    Profit = $ _____

    The profit is _______ by $ _____.
Answer #1

a) Linear Programming model is following:

Max 27P1+30P2+32P3

s.t.

1.6P1+3.1P2+2.1P3 <= 470

2.2P1+1.2P2+2.7P3 <= 370

.35P1+.35P2+.35P3 <= 70

P1, P2, P3 >= 0

b) Solution of the linear program using Excel Solver is following:

fx SUMPRODUCT(B3:D3,$B$9:$D$9) E3 A C E E G J K M N P3 1 P1 P2 S Solver Parameters X 2 Profit 32 3 27 30 5844 Set Objective:

FORMULAS:

E3 =SUMPRODUCT(B3:D3,$B$9:$D$9) copy to E5:E7

P1 = 76

P2 = 88

P3 = 36

Profit = $ 5,844

c) Total profit contribution after taking into account the setup costs = 5844 - (420+570+620) = $ 4,234

d)

Mixed-integer programming model is following:

Max 27P1+30P2+32P3-420y1-570y2-620y3

s.t.

1.6P1+3.1P2+2.1P3 <= 470

2.2P1+1.2P2+2.7P3 <= 370

.35P1+.35P2+.35P3 <= 70

P1-185y1 <= 0

P2-160y2 <= 0

P3-150y3 <= 0

P1, P2, P3 >= 0

y1, y2, y3 = 0,1

e) Solution using Excel Solver is following:

fx =SUMPRODUCT(B3:G3, $B$12:$G$12) H3 N P A C G K M O R P1 P2 1 P3 V2 y3 S Solver Parameters X 2 Profit 27 30 32 -420 -570 3

FORMULAS:

H3 =SUMPRODUCT(B3:G3,$B$12:$G$12) copy to H5:H10

P1 = 100

P2 = 100

P3 = 0

Profit = $ 4,710

The profit is increased by = 4710-4234 = $ 476